What is the molality of a solution containing 57.2 g silver perchlorate dissolved in 885 g of benzene?
1
Expert's answer
2020-04-18T07:08:55-0400

M (AgClO4) = 207.32 g/mol

n = m/M

n = 57.2 / 207.32 = 0,2759 mol




Molality = 0,2759 mol / 0.885 kg = 0.31 mol/kg.